# Generating Verifiable Chain of Thoughts from Exection-Traces

### v2

@@ 6. Conclusion and Future Work
Appendix Sample: Bi-Directional Conversation (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function sorts a list of DNA sequences by their GC-content in descending order, using lexicographical order as a tiebreaker, and includes error handling for invalid DNA sequences containing characters other than ’A’, ’T’, ’C’, ’G’. Here’s the code: def solution(dna_sequences): # Calculate GC-content using list comprehension def gc_content(seq): return (seq.count(’G’) + seq.count(’C’)) / len(seq) # Validate sequences and prepare for sorting valid_sequences = [ (gc_content(seq), seq) for seq in dna_sequences if all(nucleotide in ’ATCG’ for nucleotide in seq) ] # Raise error if any invalid sequence is found if len(valid_sequences) != len(dna_sequences): raise ValueError("Invalid DNA sequence detected.") # Sort sequences by GC-content and lexicographically valid_sequences.sort(key=lambda x: (-x[0], x[1])) # Return sorted sequences return [seq for _, seq in valid_sequences] When we pass a list of DNA sequences where each sequence is composed of a varying number of ’G’s followed by ’A’s, does the function return the sequences sorted by their GC-content from lowest to highest?
Assistant Response ### Understand The task is to determine if the given Python function solution sorts a list of DNA sequences by their GC-content from lowest to highest. Each sequence is composed of a varying number of Gs followed by As. The function calculates the GC-content for each sequence, validates the sequences, and sorts them based on GC-content and lexicographically. Key components include the gc_content function, the validation of sequences, and the sorting mechanism. ### Plan 1. **Analyze the gc_content function**: Understand how it calculates the GC-content for a sequence. 2. **Examine sequence validation**: Ensure all sequences are valid DNA sequences. 3. **Understand the sorting logic**: Determine how sequences are sorted by GC-content and lexicographically. 4. **Evaluate the overall effect**: Assess if the function sorts sequences from lowest to highest GC-content. 5. **Consider edge cases**: Check if sequences with the same GC-content are handled correctly. ### Execute 1. **GC-content Calculation**: - The gc_content function calculates the GC-content as the ratio of Gs and Cs to the total length of the sequence. Since sequences only contain Gs and As, Cs are absent, and the GC-content is simply the proportion of Gs. 2. **Sequence Validation**: - The function checks if each sequence contains only A’, T’, C’, and G’. Given the problem constraints, all sequences are valid. 3. **Sorting Logic**: - The sequences are sorted using valid_sequences.sort(key=lambda x: (-x[0], x[1]))‘. This sorts primarily by GC-content in descending order (‘-x[0]‘) and secondarily by lexicographical order (‘x[1]‘). 4. **Effect of Sorting**: - The sorting order is from highest to lowest GC-content due to the negative sign in ‘-x[0]‘. This means sequences with more Gs (higher GC-content) appear first. 5. **Edge Cases**: - If sequences have the same GC-content, they are sorted lexicographically. Since G < A’, sequences with more Gs will appear first among those with the same GC-content. ### Reflect The function does not sort sequences from lowest to highest GC-content; it sorts them from highest to lowest. The use of ‘-x[0]‘ in the sorting key indicates descending order. This conclusion aligns with the functions logic and the observed behavior in the execution trace. The reasoning process considered the functions structure, calculations, and sorting mechanism, ensuring a thorough understanding. Predicted Output: No, the function returns the sequences sorted by their GC-content from highest to lowest.
User Prompt (Backward) What kind of input list would result in the function returning a list of DNA sequences sorted from those with 0
Assistant Response ### Understand The goal is to determine what kind of input list would cause the function solution to return a list of DNA sequences sorted from those with 0% GC-content to those with 100% GC-content. The function sorts them primarily by GC-content in descending order. The final output is a list of sequences sorted in descending order of GC-content, which is the opposite of the desired order (ascending). Therefore, we need to reverse this logic to deduce the input that would result in the desired output. ### Plan 1. **Reverse the Sorting Logic**: Since the function sorts sequences by descending GC-content, we need to consider how the input could be structured so that when sorted in descending order, the sequences appear in ascending order of GC-content. 2. **Identify GC-content Calculation**: Understand how GC-content is calculated and how it influences sorting. 3. **Consider Edge Cases**: Explore sequences with extreme GC-content values (0% and 100%) and how they affect sorting. ### Execute 1. **Reverse the Sorting Logic**: - The function sorts sequences by ‘(-x[0], x[1])‘, meaning it first sorts by GC-content in descending order. To achieve ascending order of GC-content, the input should be structured such that the descending sort results in the desired order. 2. **Identify GC-content Calculation**: - GC-content is ‘(seq.count(’G’) + seq.count(’C’)) / len(seq)‘. Sequences with more G and C have higher GC-content. 3. **Deduce Input**: - To achieve the desired output, input sequences should be structured such that when sorted by descending GC-content, they naturally appear in ascending order. - Input: ‘[’AAAA’, AACG’, ACGT’, CGCG’, GGGG’]‘ - This input results in GC-content values of 0%, 50%, 50%, 100%, and 100%, respectively. When sorted by descending GC-content, the order is ‘[’GGGG’, CGCG’, ACGT’, AACG’, AAAA’]‘, which is the reverse of the desired order. ### Reflect The deduced input is structured to ensure that when sorted by descending GC-content, the sequences appear in ascending order of GC-content. This approach considers the functions sorting logic and the calculation of GC-content. Predicted Input: Plausible input 1: [’AAAA’, ’AACG’, ’ACGT’, ’CGCG’, ’GGGG’]
Appendix Sample: Bi-Directional Conversation (Difficulty 4)
User Prompt (Forward) The ‘solve_polynomial‘ function finds all real roots of a polynomial equation using Newton’s method, handling polynomials up to degree 10, with mechanisms to estimate the derivative and manage non-convergence cases. Here’s the code: def solve_polynomial(coefficients: list[float]) -> list[float]: """ Solves polynomial equations using a recursive implementation of Newton’s method. """ from math import isclose def evaluate_polynomial(coeffs, x): return sum(c * (x ** i) for i, c in enumerate(coeffs)) def evaluate_derivative(coeffs, x): return sum(i * c * (x ** (i - 1)) for i, c in enumerate(coeffs) if i != 0) def newtons_method_recursive(coeffs, x, tolerance=1e-7, max_iterations=1000, iteration=0): if iteration >= max_iterations: return None f_x = evaluate_polynomial(coeffs, x) f_prime_x = evaluate_derivative(coeffs, x) if isclose(f_prime_x, 0, abs_tol=tolerance): return None next_x = x - f_x / f_prime_x if isclose(x, next_x, abs_tol=tolerance): return next_x return newtons_method_recursive(coeffs, next_x, tolerance, max_iterations, iteration + 1) degree = len(coefficients) - 1 initial_guesses = [i * 0.5 for i in range(-degree, degree + 1)] roots = [] for guess in initial_guesses: root = newtons_method_recursive(coefficients, guess) if root is not None and all(not isclose(root, r, abs_tol=1e-5) for r in roots): roots.append(root) return roots What does the ‘solve_polynomial‘ function return when given the input ‘[1] + [0]*9 + [-1]‘?
@@
User Prompt (Backward) What input would produce the output ‘[-1.0, 1.0]‘ when passed to the ‘solve_polynomial‘ function?
[…]
Appendix Sample: Bi-Directional Conversation (Difficulty 4)
User Prompt (Forward) The ‘solution‘ function merges overlapping intervals from a list of tuples by sorting them and combining intervals that overlap or touch, ensuring the merged list is returned with a time complexity of O(n log n). Here’s the code: def solution(intervals): # Sort intervals by start time intervals.sort(key=lambda x: x[0]) # Initialize the list to store merged intervals merged_intervals = [] # Unpack the first interval and add to merged_intervals first_start, first_end = intervals[0] merged_intervals.append((first_start, first_end)) # Iterate over the remaining intervals for start, end in intervals[1:]: # Unpack the last interval in merged_intervals last_start, last_end = merged_intervals[-1] # Check for overlap if start <= last_end: # Merge intervals merged_intervals[-1] = (last_start, max(last_end, end)) else: # No overlap, add the current interval merged_intervals.append((start, end)) return merged_intervals When I run the function ‘solution‘ with the input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘, will it return ‘[(1, 10), (15, 18)]‘ as expected?
Assistant Response ### Understand The goal is to determine if the function solution correctly returns ‘[(1, 10), (15, 18)]‘ when given the input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘. The function is designed to merge overlapping intervals. It first sorts the intervals by their start times, then iteratively checks for overlaps, merging them if necessary, and finally returns the list of merged intervals. The key components influencing the reasoning are the sorting step, the loop that checks for overlaps, and the condition that determines whether intervals should be merged or added as separate intervals. ### Plan 1. **Initialization and Sorting**: Start by understanding how the input is sorted. This will affect the order in which intervals are processed. 2. **Merging Logic**: Analyze the loop that processes each interval to determine how overlaps are identified and merged. 3. **Edge Cases**: Consider edge cases such as intervals that do not overlap and intervals that are completely contained within others. 4. **Verification**: Ensure that the final list of intervals matches the expected output by following the logic step-by-step. ### Execute 1. **Initialization and Sorting**: - The input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘ is sorted by the start time, resulting in the same order since the intervals are already sorted by their start times. 2. **Merging Logic**: - Initialize merged_intervals with the first interval ‘(1, 10)‘. - Process the second interval ‘(2, 6)‘: - Compare with the last interval in merged_intervals‘, ‘(1, 10)‘. - Since ‘2 <= 10‘, they overlap. Merge them into ‘(1, 10)‘. - Process the third interval ‘(8, 10)‘: - Compare with the last interval in merged_intervals‘, ‘(1, 10)‘. - Since ‘8 <= 10‘, they overlap. Merge them into ‘(1, 10)‘. - Process the fourth interval ‘(15, 18)‘: - Compare with the last interval in merged_intervals‘, ‘(1, 10)‘. - Since ‘15 > 10‘, they do not overlap. Add ‘(15, 18)‘ as a new interval. 3. **Edge Cases**: - The intervals ‘(2, 6)‘ and ‘(8, 10)‘ are fully contained within ‘(1, 10)‘, which is correctly handled by the merging logic. - The interval ‘(15, 18)‘ does not overlap with any existing intervals, so it is added as a separate interval. 4. **Verification**: - The final list of merged intervals is ‘[(1, 10), (15, 18)]‘, which matches the expected output. ### Reflect The reasoning process confirms that the function correctly identifies and merges overlapping intervals, and correctly handles non-overlapping intervals. The expected output ‘[(1, 10), (15, 18)]‘ is indeed returned by the function when given the input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘. The approach considered both the merging logic and edge cases, ensuring a comprehensive understanding of the functions behavior. The answer is optimal and aligns with the functions intended purpose. Predicted Output: [(1, 10), (15, 18)]
User Prompt (Backward) What input would make the function ‘solution‘ return the output ‘[(1, 10), (15, 18)]‘?
Assistant Response ### Understand The goal is to determine what input to the function solution would result in the output ‘[(1, 10), (15, 18)]‘. The function takes a list of intervals and merges overlapping intervals. The process involves sorting the intervals by their start times and then iterating through them to merge any overlapping intervals. The output ‘[(1, 10), (15, 18)]‘ suggests that there are two distinct, non-overlapping intervals in the final result. This implies that any intervals that could potentially merge into a single interval must either be non-overlapping or already merged in such a way that they result in the specified output. ### Plan To deduce the input(s) that produce the output ‘[(1, 10), (15, 18)]‘, we need to reverse-engineer the merging process: 1. **Identify Non-Overlapping Intervals**: The output consists of two intervals, ‘(1, 10)‘ and ‘(15, 18)‘, which are non-overlapping. This means any input intervals that could merge into a single interval must be adjusted so they do not overlap. 2. **Consider Overlapping Scenarios**: Determine which input intervals could potentially overlap and how they must be structured to result in the given output after merging. 3. **Reverse the Merging Logic**: Work backwards from the output to infer possible inputs. This involves considering how intervals could be split or adjusted to avoid merging. 4. **Explore Multiple Configurations**: Consider different configurations of input intervals that could lead to the same output, ensuring all possibilities are covered. ### Execute 1. **Non-Overlapping Intervals**: The intervals ‘(1, 10)‘ and ‘(15, 18)‘ are already non-overlapping. Therefore, any input intervals that directly map to these without merging are valid. For example, ‘[(1, 10), (15, 18)]‘ itself is a valid input. 2. **Potential Overlaps**: - **Interval ‘(1, 10)‘**: Any input intervals that could merge into ‘(1, 10)‘ must overlap or be contiguous within this range. Possible configurations include: - ‘[(1, 10)]‘ (no merging needed) - ‘[(1, 5), (5, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 9), (9, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 6), (6, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 8), (8, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(1, 7), (7, 10)]‘ (overlapping and merge to ‘(1, 10)‘) - ‘[(2, 10), (1, 3)]‘ (overlapping and merge to ‘(1, 10)‘) 3. **Interval ‘(15, 18)‘**: Similarly, any input intervals that could merge into ‘(15, 18)‘ must overlap or be contiguous within this range. Possible configurations include: - ‘[(15, 18)]‘ (no merging needed) - ‘[(15, 17), (17, 18)]‘ (overlapping and merge to ‘(15, 18)‘) - ‘[(15, 16), (16, 18)]‘ (overlapping and merge to ‘(15, 18)‘) - ‘[(15, 19), (15, 16)]‘ (overlapping and merge to ‘(15, 18)‘) 4. **Combining Configurations**: Combine the configurations for ‘(1, 10)‘ and ‘(15, 18)‘ to form complete input sets: - ‘[(1, 10), (15, 18)]‘ - ‘[(1, 5), (5, 10), (15, 18)]‘ - ‘[(1, 9), (9, 10), (15, 18)]‘ - ‘[(1, 6), (6, 10), (15, 18)]‘ - ‘[(1, 8), (8, 10), (15, 18)]‘ - ‘[(1, 7), (7, 10), (15, 18)]‘ - ‘[(2, 10), (1, 3), (15, 18)]‘ - ‘[(1, 10), (15, 17), (17, 18)]‘ - ‘[(1, 10), (15, 16), (16, 18)]‘ - ‘[(1, 10), (15, 19), (15, 16)]‘ ### Reflect The deduced inputs cover all possible configurations that could result in the output ‘[(1, 10), (15, 18)]‘. Each configuration respects the merging logic of the function, ensuring that intervals either merge correctly or remain non-overlapping. The given input ‘[(1, 10), (2, 6), (8, 10), (15, 18)]‘ is one such configuration that results in the desired output, confirming the validity of the reasoning process. The approach was exhaustive, considering all potential overlaps and non-overlapping scenarios, ensuring no possible input was overlooked. Predicted Input: Plausible input 1: [(1, 10), (15, 18)] Plausible input 2: [(1, 5), (5, 10), (15, 18)] Plausible input 3: [(1, 9), (9, 10), (15, 18)] Plausible input 4: [(1, 6), (6, 10), (15, 18)] Plausible input 5: [(1, 8), (8, 10), (15, 18)] Plausible input 6: [(1, 7), (7, 10), (15, 18)] Plausible input 7: [(2, 10), (1, 3), (15, 18)] Plausible input 8: [(1, 10), (15, 17), (17, 18)] Plausible input 9: [(1, 10), (15, 16), (16, 18)] Plausible input 10: [(1, 10), (15, 19), (15, 16)]
Appendix Sample: Mapping an Execution Trace to Bi-Directional CoT
1. Ground-Truth Execution Trace Starting var:.. numbers = [-2, 3, -1, 0]
@@
2. Forward Reasoning constructed from Execution-Trace
@@
3. Backward Reasoning constructed from Execution-Trace
@@
What input would produce the output ‘[-2, -1, 0, 3]‘ when using the ‘parallel_sort‘ function?
Assistant Response ### Understand The goal is to determine the input(s) that would produce the output ‘[-2, -1, 0, 3]‘ when using the parallel_sort function. The function takes a list of integers and a number of threads as input. It divides the list into chunks, sorts each chunk in parallel using threads, merges the sorted chunks, and returns the fully sorted list. Key operations include list division, parallel sorting, and merging. ### Plan To deduce the input(s), we need to reverse the operations performed by the function: 1. **Reverse the Final Sort**: The output is already sorted, so the merged list before the final sort must be ‘[-2, -1, 0, 3]‘. 2. **Reverse the Merge**: Determine possible sorted chunks that, when merged, result in ‘[-2, -1, 0, 3]‘. 3. **Reverse the Chunk Sorting**: Identify possible original chunks before sorting that could result in the sorted chunks. 4. **Reverse the Chunk Division**: Determine the original list that could be divided into the identified chunks. 5. **Consider Multiple Threads**: Explore different numbers of threads to see how they affect chunk sizes and division. ### Execute 1. **Reverse the Final Sort**: The list before the final sort is ‘[-2, -1, 0, 3]‘. 2. **Reverse the Merge**: - Possible sorted chunks that merge to ‘[-2, -1, 0, 3]‘ include: - ‘[-2]‘ and ‘[-1, 0, 3]‘ - ‘[-2, -1]‘ and ‘[0, 3]‘ - ‘[-2, -1, 0]‘ and ‘[3]‘ - ‘[-2, -1, 0, 3]‘ (single chunk) 3. **Reverse the Chunk Sorting**: - For ‘[-2]‘ and ‘[-1, 0, 3]‘, possible original chunks could be ‘[-2]‘ and ‘[-1, 3, 0]‘. - For ‘[-2, -1]‘ and ‘[0, 3]‘, possible original chunks could be ‘[-2, -1]‘ and ‘[3, 0]‘. 4. **Reverse the Chunk Division**: - Consider different numbers of threads: - **2 Threads**: - Possible original lists: ‘[-2, -1, 3, 0]‘, ‘[-2, 3, -1, 0]‘, ‘[-2, -1, 0, 3]‘ - **1 Thread**: - Possible original list: ‘[-2, -1, 0, 3]‘ 5. **Verify and Explore Alternatives**: - Check if other numbers of threads (e.g., 3 or 4) could yield different valid inputs. ### Reflect The deduced inputs are consistent with the functions logic and the given output. The functions behavior with different numbers of threads and chunk divisions has been thoroughly explored. The inputs ‘[-2, -1, 3, 0]‘, ‘[-2, 3, -1, 0]‘, and ‘[-2, -1, 0, 3]‘ with 2 threads are valid, as are ‘[-2, -1, 0, 3]‘ with 1 thread, and variations with 3 or 4 threads. These inputs align with the given input ‘[-2, 3, -1, 0]‘ with 2 threads, confirming the correctness of the reasoning. Predicted Input: Plausible input 1: [-2, -1, 3, 0], 2 Plausible input 2: [-2, 3, -1, 0], 2 Plausible input 3: [-2, -1, 0, 3], 1 Plausible input 4: [-2, -1, 0, 3], 3 Plausible input 5: [-2, 3, -1, 0], 3 Plausible input 6: [-2, -1, 0, 3], 4

### v3

@@ 6 Conclusion
  • Interpretation: High agreement across independent implementations suggests correctly generated code and valid test cases (CodeT finding)
Cluster 2 (Low Consensus - REJECTED)
Solutions in this cluster: 1 out of 5 generated implementations Tests passing for this solution: 5 out of 25 generated tests
Solution in Cluster 2 Solution (Incorrect - Off-by-one error): def solution(a, b): while b > 0: a, b = b, a % b return a + 1 # Bug Tests This Solution Passes (5 out of 25): def test_zero_input(): assert solution(0, 5) == 5 def test_same_values(): assert solution(7, 7) == 7 def test_one_input(): assert solution(1, 1) == 1 def test_power_of_two(): assert solution(8, 4) == 4 def test_coprime_small(): assert solution(3, 2) == 1
Cluster 2 Metrics:
  • Solution Disagreement: 1 solution with unique behavior (no agreement)
  • Low Test Consensus: Only 5 tests, with failures indicating errors
@@
  • Interpretation: Low agreement suggests generation errors in either solutions or tests
Dual Agreement Selection:
Given 5 solutions and 25 tests (125 pairs executed), clustering reveals multiple solution groups. Cluster 1 achieves score 72.0 while Cluster 2 achieves score 5.0. Following the CodeT methodology, the highest-scoring cluster is selected, as it contains the most correctly implemented solutions and valid test cases.
Selected for Stage C:
  • Canonical Solution: Solution 1 from Cluster 1 (shortest, most readable)
  • Test Case: Selected from 18 passing tests based on coverage of the canonical solution (see below)
Test Case Selection. The highest-scoring cluster contains multiple solutions and tests (here 4 solutions × 18 tests = 72 possible pairs). To avoid redundant training samples for the same coding task, we select one (solution, test) pair per problem. We pick the canonical solution and then rank the passing tests by estimated code coverage on that solution—favoring tests whose inputs exercise more branches, loops, and edge cases, while penalizing trivial or repeated inputs. We compared this coverage-based selection against an LLM-as-judge approach and found comparable downstream performance, so we use the simpler coverage-based criterion.
This pipeline demonstrates Dual Agreement’s effectiveness: by clustering solutions based on test consensus, it identifies the highest-quality code-test pairs, and coverage-based selection ensures each problem contributes one representative, high-coverage training sample to Stage C.
CoT Samples: Mapping an Execution Trace to Bi-Directional CoT
We present representative examples of verifiable CoT data generated by our pipeline, covering forward, backward, and bi-directional reasoning on hard (Difficulty 4) problems. Each sample shows the user prompt and the model’s trace-grounded response.
This example walks through our complete pipeline on a single Difficulty 4 training sample involving multi-threaded parallel sorting. We show the problem instruction, canonical solution selected via Dual Agreement, test case with maximum coverage, pysnooper instrumentation, and the resulting execution trace that grounds the bi-directional CoT.
Pipeline Inputs for Trace Generation Problem Instruction (Difficulty 4): Implement a function parallel_sort that takes a list of integers and a number of threads, divides the list into chunks, sorts each chunk in a separate thread, and merges the sorted chunks into a single sorted list. Canonical Solution (selected by Dual Agreement): import threading def parallel_sort(numbers: list[int], num_threads: int) -> list[int]: if not numbers: return [] chunk_size = len(numbers) // num_threads chunks = [numbers[i:i + chunk_size] for i in range(0, len(numbers), chunk_size)] threads = [] sorted_chunks = [None] * num_threads lock = threading.Lock() def sort_chunk(index): sorted_chunks[index] = sorted(chunks[index]) with lock: print(f"Chunk {index} sorted: {sorted_chunks[index]}") for i in range(num_threads): thread = threading.Thread(target=sort_chunk, args=(i,)) threads.append(thread) thread.start() for thread in threads: thread.join() merged_list = [] for chunk in sorted_chunks: if chunk: merged_list.extend(chunk) return sorted(merged_list) Test Case (maximum coverage): assert parallel_sort([-2, 3, -1, 0], 2) == [-2, -1, 0, 3] Pysnooper Instrumentation: We instrument the canonical solution and run it with the test case: import pysnooper @pysnooper.snoop() def parallel_sort(numbers, num_threads): ... # same function body as above # Run with the test case parallel_sort([-2, 3, -1, 0], 2) The execution trace below is the resulting pysnooper output.
Raw Execution Trace (pysnooper output) Starting var:.. numbers = [-2, 3, -1, 0]
@@
From the execution trace above, we construct training samples in a chat-style format with alternating User and Assistant turns. The three reasoning directions differ in how these turns are arranged:
  • Forward-only: User (instruction + code + input question) $\rightarrow$ Assistant (CoT + predicted output)
  • Backward-only: User (instruction + code + output question) $\rightarrow$ Assistant (CoT + predicted input)
  • Bi-directional: User (forward question) $\rightarrow$ Assistant (forward CoT) $\rightarrow$ User (backward question) $\rightarrow$ Assistant (backward CoT)
Below, we show the forward and backward turns for this example. A forward-only sample uses only the first turn; a backward-only sample uses only the second; a bi-directional sample concatenates both.
@@
What input (list and number of threads) to ‘parallel_sort‘ produces the output ‘[-2, -1, 0, 3]‘?
Assistant Response ### Understand The function parallel_sort takes a list of integers and a number of threads. It divides the list into chunks, sorts each chunk in a separate thread, merges the sorted chunks into merged_list‘, and returns sorted(merged_list)‘. The output is ‘[-2, -1, 0, 3]‘. To find the input, I need to work backward through each operation. ### Plan 1. Start from the output ‘[-2, -1, 0, 3]‘ and reverse the final sorted()‘ call. 2. Reverse the merge to find the sorted chunks. 3. Reverse the per-chunk sorting to find the original chunks. 4. Reverse the chunk division to recover the input list and number of threads. ### Execute 1. **Reverse the final sort**: The function returns sorted(merged_list) = [-2, -1, 0, 3]‘. The merged_list before sorting must contain the same elements. Since merging extends chunks sequentially, merged_list was built by appending ‘[-2, 3]‘ followed by ‘[-1, 0]‘, giving merged_list = [-2, 3, -1, 0]‘. Applying sorted()‘ to this yields ‘[-2, -1, 0, 3]‘, which matches. 2. **Reverse the merge**: The merge loop iterates over sorted_chunks and extends merged_list‘. Since merged_list = [-2, 3, -1, 0]‘ and the first extension added ‘[-2, 3]‘ and the second added ‘[-1, 0]‘, we get sorted_chunks = [[-2, 3], [-1, 0]]‘. 3. **Reverse the chunk sorting**: Each chunk was sorted individually by sort_chunk‘. For the first chunk, sorted(chunk_0) = [-2, 3]‘. Since ‘[-2, 3]‘ is already sorted, the original chunk_0 = [-2, 3]‘. For the second chunk, sorted(chunk_1) = [-1, 0]‘. Since ‘[-1, 0]‘ is already sorted, the original chunk_1 = [-1, 0]‘. 4. **Reverse the chunk division**: The chunks ‘[[-2, 3], [-1, 0]]‘ have size 2 each. chunk_size = len(numbers) // num_threads‘, so ‘2 = 4 // num_threads‘, giving num_threads = 2‘. The original list is the concatenation of the chunks: ‘[-2, 3] + [-1, 0] = [-2, 3, -1, 0]‘. ### Reflect By reversing each stage---final sort, merge, chunk sorting, and chunk division---we recover the full input. The deduced numbers = [-2, 3, -1, 0]‘ with num_threads = 2‘ produces chunks ‘[[-2, 3], [-1, 0]]‘, which sort to ‘[[-2, 3], [-1, 0]]‘, merge to ‘[-2, 3, -1, 0]‘, and sorted([-2, 3, -1, 0])‘ returns ‘[-2, -1, 0, 3]‘, confirming consistency. Predicted Input: [-2, 3, -1, 0], 2
Sliding Window-based CoT Verification
Since natural language explanations do not map one-to-one to trace lines (e.g., one sentence may summarize several steps), we use a sliding window algorithm to verify each rationale against the trace. The verifier maintains a pointer to the current position in the execution trace. For each CoT sentence, we extract verifiable entities—variable names, values, and control flow keywords—using an LLM (extraction only, not reasoning). For each extracted entity, the algorithm scans a lookahead window of K=15K{=}15 trace steps from the current pointer. A reasoning step is accepted if:
  1. 1.
    Event Matching: A trace step within the window contains the cited variable modification or value.
  2. 2.
    State Consistency: The cited value exists in the program’s accumulated state at the current pointer.
If neither condition holds, the entity is ungrounded and the rationale is flagged. As a second check, the final answer stated in the rationale is string-matched against ground-truth I/O from the test case. Rationales failing either check are discarded. This process rejected approximately 30% of generated rationales.
Detailed Rejection Example
Below we show a full trace and rejected rationale side by side for binary_search([1,3,5,7], 5), followed by the verification walkthrough.
Full Trace [1] Starting var: arr=[1,3,5,7], target=5
[2] line lo = 0
[3] New var: lo = 0
[4] line hi = len(arr) - 1
[5] New var: hi = 3
[6] line while lo <= hi:
[7] line mid = (0+3)//2
[8] New var: mid = 1
[9] line if arr[1]==5 $\to$ False
[10] line elif arr[1]<5 $\to$ True
[11] line lo = mid + 1
[12] Modified var: lo = 2
[13] line while lo <= hi:
[14] line mid = (2+3)//2
[15] Modified var: mid = 2
[16] line if arr[2]==5 $\to$ True
[17] line return mid
[18] Return value: 2
Rejected Rationale ### Understand The function performs binary search on a sorted array to find the index of a target value. ### Plan 1. Initialize lo=0, hi=len(arr)-1. 2. Compute mid, compare arr[mid] with target. 3. Narrow search range until found or exhausted. ### Execute 1. Start with lo=0 and hi=3. 2. First iteration: mid=(0+3)//2=1. arr[1]=3, which is less than target 5, so lo becomes 2. 3. Second iteration: mid=(2+3)//2=2. arr[2]=5. Since 5>5 is false and 5<5 is false, we enter the else branch and set hi=1. 4. Loop continues but lo>hi, so we exit. ### Reflect The search narrowed correctly in the first iteration but failed to find the target in the second, returning -1. Predicted Output: -1
Verification Walkthrough Sentence 1: “We start with lo=0 and hi=3.”
Extracted entities: lo=0, hi=3
Window scan (pointer at [1]): lo=0 found at [3] ✓, hi=3 found at [5] ✓. Pointer advances to [5].
Sentence 2: “mid = (0+3)//2 = 1. arr[1]=3 < 5, so lo becomes 2.”
Extracted entities: mid=1, arr[1]=3, lo=2
Window scan: mid=1 found at [8] ✓, arr[1]<target matches [10] ✓, lo=2 found at [12] ✓. Pointer advances to [12].
Sentence 3: “mid = (2+3)//2 = 2. arr[2]=5.”
Extracted entities: mid=2, arr[2]=5
Window scan: mid=2 found at [15] ✓, arr[2]==5 matches [16] ✓. Pointer advances to [16].
Sentence 4: “Since 5>5 is false and 5<5 is false, we enter the else branch and set hi=1.”
Extracted entities: else branch, hi=1
Window scan ([16]–[18]+$K$): Trace shows if arr[2]==5 $\to$ True at [16], followed by return mid at [17]. No else branch found. hi=1 not found in window or accumulated state. Event Matching fails.
Sentence 5: “Return -1.”
Extracted entities: return=-1
Trace shows Return value: 2 at [18].
I/O check also fails.